OpenStudy (anonymous):

cotx = -2 and secx < 0 How do I find sinx, cosx, and tanx?

OpenStudy (anonymous):

r u given any more info?

jhonyy9 (jhonyy9):

so you know that cot x = cos x / sin x and tan x = 1/cot x yes ? hope this will help you

OpenStudy (guptalakshay558):

Given :- Cotx = -2 secx<0 Approach:- We know cotx and secx are both together negative in only 2nd quadrant Solution :- cotx =-2 so, tanx = -1/2 (cotx = 1/tanx) sinx = \[1/\sqrt{5}\]( sinx is positive in second quadrant) cos= \[-2/\sqrt{5}\] (cos x is negative in seond quadrent )